OpenStudy (anonymous):

what does it mean to find the area of the shaded region? The shaded area is to the right of z=2.04

OpenStudy (anonymous):

What shape is the shaded region? Do you have a picture?

OpenStudy (anonymous):

It's the shape of a bell. And on the right hand side of the tail is shaded.

OpenStudy (anonymous):

Do you have the equation of the line?

OpenStudy (anonymous):

The problem says to draw a standard normal curve and shade the area to the right of z= 2.04. Then find the area of the shaded region.

OpenStudy (anonymous):

Thank you the book shows that the answer is 0.0207 but I just don't understand how they get that. I know when I look under the Standard Normal Distribution chart that answer is under -2.0 but why wouldn't I look under 2.0 instead of -?

OpenStudy (zarkon):

\[P(Z>2.04)=1-P(Z\le 2.04)\] or by the symmetry of the standard normal distributuon \[P(Z>2.04)=P(Z<-2.04)\]

OpenStudy (anonymous):

Thank you!